Find the length of the curve given by x=t^3, y=2t^2 in 0<=t<=1. What is the slope of the curve at t=1/2.
1
Expert's answer
2020-05-12T18:31:14-0400

Given equation of of the curve in parametric form is "x=t^3, \: y=2t^2" . Now, we have to find the length of this curve from "0 \leq t \leq 1" .

Since, length of a parametric curve is given by,

"L=\int_{t_1}^{t_2} \sqrt{(\frac{dx}{dt})^2 + (\frac{dy}{dt})^2} dt \hspace{1.5cm} (\star)"


Thus, plugin

"\frac{dx}{dt}=3t^2 \: \& \frac{dy}{dt}=4t"

in equation "(\star)", we get

"L= \int_{0}^{1} \sqrt{(3t^2)^2 + (4t)^2} dt"

"\implies L= \int_{0}^{1} t\sqrt{9t^2 + 16} dt"


Let, substitute "u= 9t^2 +16 \implies \frac{du}{18t}=dt" and "\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 u & 16 & 25\\ \hline
 t & 0 & 1 \\
\end{array}" , hence

"L=\frac{1}{18}\int_{16}^{25} \sqrt{u} \:du\\
L=\frac{1}{18}\frac{ u^{\frac{3}{2}}}{\frac{3}{2}} \bigg|_{16}^{25} \\L=\frac{1}{27}u^{\frac{3}{2}}\bigg|_{16}^{25}\\L=\frac{1}{27}(25^{\frac{3}{2}}-16^{\frac{3}{2}})=\frac{61}{27}"

Therefore, the required length is "\frac{61}{27}" .

Now we know that if "y=f(x)" is a parametric curve such that "x=x(t) \&y=y(t)" ,then slope of that curve at a point "P" with given parameter "t=t_0" is given by,

"y'=\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\bigg|_{t=t_0}\hspace{1.5cm} (\spades)"

Thus, slope of the given curve at "t=\frac{1}{2}" is given by "\frac{dy}{dx}" ,thus from equation"(\spades)" we get,


"\frac{dy}{dx}=\frac{4}{3t} |_{t=\frac{1}{2}}=\frac{8}{3}"

Hence, we are done.